OpenStudy (anonymous):

Express the following equation in slope-intercept form: 17x - 10y = 80

OpenStudy (anonymous):

-10y=-17x+80 y= 1.7-8

OpenStudy (anonymous):

y=1.7x-.8

OpenStudy (anonymous):

y=\[\frac{ 17 }{ 10 }x-\frac{ 8 }{ 10 }\]